OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, –3) and has a slope of 2? Answer a.2x + y = –5 b.2x – y = –5 c.2x – y = 5 d.2x + y = 5

OpenStudy (anonymous):

c. 2x-y=5

OpenStudy (anonymous):

-3 - y= 2( 1 - x ) -3 -y = 2 -2x 2x-y= 5 = c.